# A simpler and parallelizable $O(\sqrt{\log n})$-approximation algorithm for Sparsest Cut

### v4

@@ 1 Introduction
The line of works above culminated in the result of Sherman [She09], who showed how to compute $O(\sqrt{(\log n)/\varepsilon})$-approximation for Sparsest Cut using $O(n^{\varepsilon}\log^{O(1)}n)$ maxflows for any $\varepsilon\in[\Theta(1/\log n),\Theta(1)]$. This effectively subsumes previous results, as taking $\varepsilon=\Theta(1/\log n)$ yields an $O(\log n)$ approximation using $O(\log^{O(1)}n)$ maxflows, while a sufficiently small constant $\varepsilon$ achieves an $O(\sqrt{\log n})$-approximation and improves on the $\tilde{O}(n^{2})$ runtime in [AHK10, AK16]. In particular, using the recent almost linear-time maxflow algorithm [${}^{+}$22] yields $O(n^{1+\varepsilon})$ complexity for $O(\sqrt{\log n})$-approximation. (As usual, we assume in this paper that graph $G$ has $m=O(n\log n)$ edges. This can be achieved by sparsifying the graph using the algorithm of Benczúr and Karger [BK96], which with high probability preserves the cost of all cuts up to any given constant factor.)
@@
(ii) The algorithm is parallelizable: it can be implemented on $O(n^{\varepsilon})$ processors in expected parallel runtime $O((\log^{O(1)}n)\cdot T_{\tt maxflow})$ (in any version of the PRAM model), where “parallel runtime” is defined as the maximum runtime over all processors. Note that this is an exponential improvement over complexity $O((n^{\varepsilon}\log^{O(1)}n)\cdot T_{\tt maxflow})$ of the sequential version.
@@
In order to compute violating paths, we first design procedure ${\tt Matching}(u)$ that takes vector $u\in\mathbb{R}^{d}$ and outputs a directed matching on a given set $S\subset\mathbb{R}^{d}$ with $|S|=\Theta(n)$. (It works by calling a maxflow algorithm and then postprocessing the flow). The problem then boils down to the following: given vector $u$ randomly sampled from the Gaussian distribution, we need to sample vectors $u_{1},\ldots,u_{K}$ so that set ${\tt Matching}(u_{1})\circ\ldots\circ{\tt Matching}(u_{K})$ contains many paths $(x_{0},x_{1},\ldots,x_{K})$ with “large stretch”, i.e. with $\langle x_{K}-x_{0},u\rangle\geq\Omega(K)$. (Here $\circ$” is the operation that “chains together” paths in a natural way). This was also a key task in [She09], where it was needed for implementing one inner MW step.
@@
Another very recent paper by Agarwal-Khanna-Li-Patil-Wang-White-Zhong [${}^{+}$24] presented a parallel algorithm for approximate maxflow with polylogarithmic depth and near-linear work in the PRAM model. Using this algorithm, they presented, in particular, an $O(\log^{3}n)$-approximation sparsest cut algorithm with polylogarithmic depth and near-linear work (by building on the work [NS17] who showed how to replace exact maxflow computations in the cut-matching game [KRV06, OSVV08] with approximate maxflows).
@@ 2 Background: Arora-Kale framework
Let $c_{e}\geq 0$ be the weight of edge $e$ in $G$. The standard SDP relaxation of the $c$-Balanced Separator problem can be written in vector form as follows:
@@
The optimum of this SDP divided by 4 is a lower bound on the minimum $c$-Balanced Separator problem. Arora and Kale considered a slightly different relaxation:
@@
Here $p$ stands for a path $p=(p_{0},\ldots,p_{\ell(p)})$ in graph $G$, notation “$\forall S$” means all subsets $S\subseteq V$ of size at least $(1-c/4)n$, and $a=3c-4c^{2}$. Note that triangle inequalities (1c) imply path inequalities (2c), while constraints (1b) and (1d) imply constraints (2d) (see [AK16]). SDP (2) may be looser than (2), but its optimum divided by 4 is still a lower bound on the minimum $c$-Balanced Separator problem.
@@ 2.1 Matrix multiplicative weights algorithm
To solve the above SDP, [AK16] converts an optimization problem to a feasibility problem by replacing the objective (2a) with the constraint
$C\bullet X\leq\alpha$ (${}^{\prime}$)
@@
Input: vectors $v_{1},\ldots,v_{n}\in\mathbb{R}^{n}$ and $\tilde{v}_{1},\ldots,\tilde{v}_{n}\in\mathbb{R}^{d}$ satisfying $||\tilde{v}_{i}||^{2}$   $\;\;\leq\;\;2$ $\forall i$ (4a) $\sum_{i,j\in V:i<j}||\tilde{v}_{i}-\tilde{v}_{j}||^{2}$ $\;\;\geq\;\;\tfrac{an^{2}}{4}$ (4b) $|\;||\tilde{v}_{i}||^{2}-||v_{i}||^{2}\;|$ $\;\;\leq\;\;\gamma(||\tilde{v}_{i}||^{2}+\tau)$ $\forall i$ (4c) $|\;||\tilde{v}_{i}-\tilde{v}_{j}||^{2}-||v_{i}-v_{j}||^{2}\;|$ $\;\;\leq\;\;\gamma(||\tilde{v}_{i}-\tilde{v}_{j}||^{2}+\tau)$ $\forall i,j$ (4d) for some constants $\gamma,\tau>0$. Let $X=V^{T}V$ and $\tilde{X}=\tilde{V}^{T}\tilde{V}$ where $V\in\mathbb{R}^{n\times n}$ and $\tilde{V}\in\mathbb{R}^{d\times n}$ are the matrices with columns $\{v_{i}\}$ and $\{\tilde{v}_{i}\}$, respectively. Output: either (i) variables $f_{p}\geq 0$ and symmetric matrix $F\preceq C$ such that $\left(\sum\nolimits_{p}f_{p}T_{p}-F\right)\bullet X\leq-\alpha$ (5) or (ii) a $\Theta(1)$-balanced cut of value at most $\kappa\alpha$.
The number of iterations of the MW algorithm will depend on the maximum possible spectral norm of matrix $\tfrac{\alpha}{n}I+\sum_{p}f_{p}T_{p}-F$. This parameter is called the width of the oracle, and will be denoted as $\rho$. We will use the bound $\rho=||\tfrac{\alpha}{n}I+\sum_{p}f_{p}T_{p}-F||\leq\tfrac{\alpha}{n}+||\sum_{p}f_{p}T_{p}-F||$.
@@ 3 Our algorithm
Note that we do not require paths in $M$ to be paths in the original graph $G$. Thus, we need to use SDP relaxation (2) with all possible paths $p$ (with a certain bound on the length), which is still a valid relaxation of the problem.
@@
    In contrast, parallelizing an MW algorithm cay be a difficult task. (There exists some work on parallelizing algorithms involving MW, e.g. [CQ19, ASZ20]; however, they consider very specific problems.)
@@ 3.1 Correlated Gaussians and measure concentration
We write $u\sim{\cal N}$ to indicate that $u$ is a random vector in $\mathbb{R}^{d}$ with Gaussian independent components $u_{i}\sim{\cal N}(0,1)$. Throughout the paper notation ${\tt Pr}_{u}[\cdot]$ means the probability under distribution $u\sim{\cal N}$. We write $(u,u^{\prime})\sim{\cal N}_{\omega}$ for $\omega\in[0,1)$ to indicate that $(u,u^{\prime})$ are random vectors in $\mathbb{R}^{d}\times\mathbb{R}^{d}$ such that for each $i\in[d]$, pair $(u_{i},u^{\prime}_{i})$ is an independent 2-dimensional Gaussian with mean $(0,0)^{T}$ and covariance matrix $\begin{pmatrix}1&\omega\\ \omega&1\end{pmatrix}$. We write $u^{\prime}\sim_{\omega}u$ to indicate that $u^{\prime}$ is an $\omega$-correlated copy of $u$ [She09], i.e. $(u,u^{\prime})$ is generated according to $(u,u^{\prime})\sim{\cal N}_{\omega}$ conditioned on fixed $u$. It can be checked that for each $i\in[d]$, $u^{\prime}_{i}$ is an independent Gaussian with mean $\omega\cdot u_{i}$ and variance $1-\omega^{2}$. Note, if $(u,u^{\prime})\sim{\cal N}_{\omega}$ then $u\sim{\cal N}$ and $u^{\prime}\sim{\cal N}$. Conversely, the process $u\sim{\cal N},\hat{u}\sim_{\omega}u$ generates pair $(u,u^{\prime})$ with distribution ${\cal N}_{\omega}$. The same is true for the process $u^{\prime}\sim{\cal N},u\sim_{\omega}u^{\prime}$.
The key property for obtaining an $O(\sqrt{\log n})$-approximation algorithm is measure concentration of the Gaussian distribution. This property can be expressed in a number of different ways; we will use the following version. [^2] Theorem 5 is formulated in [${}^{+}$06] for the discrete cube, but the proof also works for the Gaussian distribution. For completeness, we reproduce the proof in Appendix B.
Theorem 5 ([${}^{+}$06]).
@@ 3.2 Procedure ${\tt Matching}(u)$
1 compute $w_{x}=\langle x,u\rangle$ for each $x\in V$
@@
For the purpose of analysis we make the following assumption: if Algorithm 1 terminates at line 5 or 7 then it returns $\varnothing$ (the empty matching). Thus, we always have ${\tt Matching}(u)\subseteq V\times V$ and $|{\tt Matching}(u)|\leq|V|$. The following result is proved in Appendix C.
@@ 3.3 Matching covers
(c) There holds ${\tt Pr}_{u}[u\mbox{ is regular\,}]\geq 1-\tfrac{1}{n}$.
@@ 3.4 Chaining algorithms
sample $(u_{1},\ldots,u_{K})\sim{\cal N}_{\omega}^{K}|u_{1}$ where $\omega=1-1/K$                                  // equivalently, sample $u_{2}\!\sim_{\omega}\!u_{1}$, $u_{3}\!\sim_{\omega}\!u_{2},~{}\ldots,~{}u_{K}\!\sim_{\omega}\!u_{K-1}$
@@
Let $x=\mathbb{E}_{u\sim{\cal N}|u\mbox{\scriptsize~{}is regular},M\sim{\cal M}(u)}[|M|]=\mathbb{E}_{u\sim{\cal N}|u\mbox{\scriptsize~{}is regular},M\sim{\cal M}(u)}[|M^{\tt violating}|]$ (by Lemma 8(b)), $y=\mathbb{E}_{u\sim{\cal N}|u\mbox{\scriptsize~{}is not regular},M\sim{\cal M}(u)}[|M|]\leq n$ and $p={\tt Pr}_{u}[u\mbox{ is not regular}]\leq\tfrac{1}{n}$ (by Lemma 8(c)). We have $n^{1-\varepsilon}\leq\mathbb{E}_{M\sim{\cal M}({\cal N})}[|M|]=(1-p)x+py\leq(1-p)x+\tfrac{1}{n}\cdot n$ and so $(1-p)x\geq n^{1-\varepsilon}-1$. Therefore, $\mathbb{E}_{M\sim{\tt SamplePaths}^{K}({\cal N})}[|M^{\tt violating}|]\geq(1-p)x\geq n^{1-\varepsilon}-1$. ∎

### v5

@@ 1 Introduction
The line of works above culminated in the result of Sherman [She09], who showed how to compute $O(\sqrt{(\log n)/\varepsilon})$-approximation for Sparsest Cut using $O(n^{\varepsilon}\log^{O(1)}n)$ maxflows for any $\varepsilon\in[\Theta(1/\log n),\Theta(1)]$. This effectively subsumes previous results, as taking $\varepsilon=\Theta(1/\log n)$ yields an $O(\log n)$ approximation using $O(\log^{O(1)}n)$ maxflows, while a sufficiently small constant $\varepsilon$ achieves an $O(\sqrt{\log n})$-approximation and improves on the $\tilde{O}(n^{2})$ runtime in [AHK10, AK16]. In particular, using the recent almost linear-time maxflow algorithm [CKL+22] yields $O(n^{1+\varepsilon})$ complexity for $O(\sqrt{\log n})$-approximation. (As usual, we assume in this paper that graph $G$ has $m=O(n\log n)$ edges. This can be achieved by sparsifying the graph using the algorithm of Benczúr and Karger [BK96], which with high probability preserves the cost of all cuts up to any given constant factor.)
@@
(ii) The algorithm is parallelizable: it can be implemented on $O(n^{\varepsilon})$ processors in expected parallel runtime $O((\log^{O(1)}n)\cdot T_{\tt maxflow})$ (in any version of the PRAM model), where “parallel runtime” is defined as the maximum runtime over all processors.
@@
In order to compute violating paths, we first design procedure ${\tt Matching}(u)$ that takes vector $u\in\mathbb{R}^{d}$ and outputs a directed matching on a given set $S\subset\mathbb{R}^{d}$ with $|S|=\Theta(n)$. (It works by calling a maxflow algorithm and then postprocessing the flow). The problem then boils down to the following: given vector $u$ randomly sampled from the Gaussian distribution, we need to sample vectors $u_{1},\ldots,u_{K}$ so that set ${\tt Matching}(u_{1})\circ\ldots\circ{\tt Matching}(u_{K})$ contains many paths $(x_{0},x_{1},\ldots,x_{K})$ with “large stretch”, i.e. with $\langle x_{K}-x_{0},u\rangle\geq\Omega(K)$. (Here $\circ$” is the operation that “chains together” paths in a natural way). This was also a key task in [She09], where it was needed for implementing one inner MW step. We refer to an algorithm that samples vectors $u_{1},\ldots,u_{K}$ as a chaining algorithm.
@@
Another very recent paper by Agarwal-Khanna-Li-Patil-Wang-White-Zhong [AKL+24] presented a parallel algorithm for approximate maxflow with polylogarithmic depth and near-linear work in the PRAM model. Using this algorithm, they presented, in particular, an $O(\log^{3}n)$-approximation sparsest cut algorithm with polylogarithmic depth and near-linear work (by building on the work [NS17] who showed how to replace exact maxflow computations in the cut-matching game [KRV06, OSVV08] with approximate maxflows).
@@ 2 Background: Arora-Kale framework
Let $c_{e}\geq 0$ be the weight of edge $e$ in $G$. The standard SDP relaxation of the $c$-Balanced Separator problem can be written in vector form as follows (see [ARV09]):
@@
Here $v_{i}\in\mathbb{R}^{n}$ for each node $i\in V$. The optimum of this SDP divided by 4 is a lower bound on the minimum $c$-Balanced Separator problem. Arora and Kale considered a slightly different relaxation:
@@
Here $p$ stands for a path $p=(p_{0},\ldots,p_{\ell(p)})$ in graph $G$, notation “$\forall S$” means all subsets $S\subseteq V$ of size at least $(1-c/4)n$, and $\xi=3c-4c^{2}$. Matrix $X$ is defined via $X=V^{T}V$ where $V$ is the $n\times n$ matrix with columns $v_{1},\ldots,v_{n}$. We have $C={\cal L}(G)$ and $T_{p}={\cal L}(G_{p})-{\cal L}(G_{(p_{0},p_{\ell(p)})})$ where ${\cal L}(\cdot)$ is the Laplacian of the corresponding graph and $G_{q}$ for a path $q$ is the undirected unweighted graph containing all edges of $q$. Finally, $K_{S}(i,i)=|S|-1$ for $i\in S$, $K_{S}(i,j)=-1$ for distinct $i,j\in S$, and $K_{S}(i,j)=0$ in all other cases.
Note that triangle inequalities (1c) imply path inequalities (2c), while constraints (1b) and (1d) imply constraints (2d) (see [AK16]). SDP (2) may be looser than (1), but its optimum divided by 4 is still a lower bound on the minimum $c$-Balanced Separator problem.
In the sequel we will use a slight modification of (2) in which constraints (2c) are enforced for all sequences $p=(p_{0},\ldots,p_{\ell(p)})$ of distinct nodes in $G$, and not just paths in $G$. This modification can only make the relaxation stronger; constraints (2c) are now equivalent to the triangle inequalities in (1c).
@@ 2.1 Matrix multiplicative weights algorithm
To solve the above SDP, [AK16] converts it to a sequence of feasibility problems. First, an interval $[L,U]$ is computed containing an optimal value of the objective. One can use, for example, the algorithm in [OSVV08] to get such an interval with $U/L=O(\log n)$ via $O(\log^{O(1)}n)$ maxflows. Fix a value $\alpha\in[L,U]$, and replace the objective (2a) with the constraint
$C\bullet X\leq\alpha$ (2a)
It suffices to try $O(\log(U/L))=O(\log\log n)$ values of threshold $\alpha$ if we are willing to accept the loss by a constant factor in the approximation ratio.
@@
Input: vectors $v_{1},\ldots,v_{n}\in\mathbb{R}^{n}$ and $\tilde{v}_{1},\ldots,\tilde{v}_{n}\in\mathbb{R}^{d}$ satisfying $||\tilde{v}_{i}||^{2}$   $\;\;\leq\;\;2$ $\forall i$ (4a) $\sum_{i,j\in V:i<j}||\tilde{v}_{i}-\tilde{v}_{j}||^{2}$ $\;\;\geq\;\;\tfrac{\xi n^{2}}{4}$ (4b) $|\;||\tilde{v}_{i}||^{2}-||v_{i}||^{2}\;|$ $\;\;\leq\;\;\gamma(||\tilde{v}_{i}||^{2}+\tau)$ $\forall i$ (4c) $|\;||\tilde{v}_{i}-\tilde{v}_{j}||^{2}-||v_{i}-v_{j}||^{2}\;|$ $\;\;\leq\;\;\gamma(||\tilde{v}_{i}-\tilde{v}_{j}||^{2}+\tau)$ $\forall i,j$ (4d) for some constants $\gamma,\tau>0$. Let $X=V^{T}V$ and $\tilde{X}=\tilde{V}^{T}\tilde{V}$ where $V\in\mathbb{R}^{n\times n}$ and $\tilde{V}\in\mathbb{R}^{d\times n}$ are the matrices with columns $\{v_{i}\}$ and $\{\tilde{v}_{i}\}$, respectively. Output: either (i) variables $f_{p}\geq 0$ and symmetric matrix $F\preceq C$ such that $\left(\sum\nolimits_{p}f_{p}T_{p}-F\right)\bullet X\leq-\alpha$ (5) or (ii) a $\Theta(1)$-balanced cut of value at most $\kappa\alpha$.
The number of iterations of the MW algorithm will depend on the maximum possible spectral norm of matrix $\tfrac{\alpha}{n}I+\sum_{p}f_{p}T_{p}-F$. This parameter is called the width of the oracle, and will be denoted as $\rho$. We will use the bound $\rho=||\tfrac{\alpha}{n}I+\sum_{p}f_{p}T_{p}-F||\leq\tfrac{\alpha}{n}+||\sum_{p}f_{p}T_{p}-F||$. The algorithm has the following guarantee (see Appendix A for details).
@@ 3 Our algorithm
Recall that in the specification of the oracle we required paths $p$ to have distinct nodes (to make the number of constraints finite). This does not cause problems for Lemma 3: if some path in $M$ does not satisfy this, then before applying the lemma we can shorten it while preserving endpoints and condition (8).
@@ 3.1 Correlated Gaussians and measure concentration
We write $u\sim{\cal N}$ to indicate that $u$ is a random vector in $\mathbb{R}^{d}$ with Gaussian independent components $u_{i}\sim{\cal N}(0,1)$. Throughout the paper notation ${\tt Pr}_{u}[\cdot]$ means the probability under distribution $u\sim{\cal N}$. We write $(u,u^{\prime})\sim{\cal N}_{\omega}$ for $\omega\in[0,1)$ to indicate that $(u,u^{\prime})$ are random vectors in $\mathbb{R}^{d}\times\mathbb{R}^{d}$ such that for each $i\in[d]$, pair $(u_{i},u^{\prime}_{i})$ is an independent 2-dimensional Gaussian with mean $(0,0)^{T}$ and covariance matrix $\begin{pmatrix}1&\omega\\ \omega&1\end{pmatrix}$. We write $u^{\prime}\sim_{\omega}u$ to indicate that $u^{\prime}$ is an $\omega$-correlated copy of $u$ [She09], i.e. $(u,u^{\prime})$ is generated according to $(u,u^{\prime})\sim{\cal N}_{\omega}$ conditioned on fixed $u$. It can be checked that for each $i\in[d]$, $u^{\prime}_{i}$ is an independent Gaussian with mean $\omega\cdot u_{i}$ and variance $1-\omega^{2}$. Note, if $(u,u^{\prime})\sim{\cal N}_{\omega}$ then $u\sim{\cal N}$ and $u^{\prime}\sim{\cal N}$. Conversely, the process $u\sim{\cal N},u^{\prime}\sim_{\omega}u$ generates pair $(u,u^{\prime})$ with distribution ${\cal N}_{\omega}$. The same is true for the process $u^{\prime}\sim{\cal N},u\sim_{\omega}u^{\prime}$.
The key property for obtaining an $O(\sqrt{\log n})$-approximation algorithm is measure concentration of the Gaussian distribution. This property can be expressed in a number of different ways; we will use the following version. [^2] Theorem 5 is formulated in [MOR+06] for the discrete cube, but the proof also works for the Gaussian distribution. For completeness, we reproduce the proof in Appendix B.
Theorem 5 ([MOR+06]).
@@ 3.2 Procedure ${\tt Matching}(u)$
1
2compute $w_{x}=\langle x,u\rangle$ for each $x\in V$
@@
For the purpose of analysis we make the following assumption: if Algorithm 1 terminates at line 5 or 7 then it returns $\varnothing$ (the empty matching). Thus, we always have ${\tt Matching}(u)\subseteq V\times V$ and $|{\tt Matching}(u)|\leq|V|$.
@@
Proof.
(b)   Assume that condition (ii) is false. By a standard argument, conditions (4a) and (4b) imply the following: there exist constants $c^{\prime}\in(0,c)$, $\beta>0$ and $\sigma>0$ such that with probability at least $\beta$ Algorithm 1 reaches line 9 and we have $w_{y}-w_{x}\geq\sigma$ for all $x\in A$, $y\in B$ (see [Kal07, Lemma 14]). Suppose that this event happens. We claim that in this case $|M|\geq\tfrac{1}{3}c^{\prime}n$. Indeed, suppose this is false. Let $A^{\prime}\subseteq A$ and $B^{\prime}\subseteq B$ be the sets of nodes involved in $M$ (with $|A^{\prime}|=|B^{\prime}|=|M|=k$). The total value of flow from $A$ to $B$ is at least $c^{\prime}\pi n$ (otherwise we would have terminated at line 5). The value of flow leaving $A^{\prime}$ is at most $|A^{\prime}|\cdot\pi\leq\tfrac{1}{3}c^{\prime}\pi n$. Similarly, the value of flow entering $B^{\prime}$ is at most $|B^{\prime}|\cdot\pi\leq\tfrac{1}{3}c^{\prime}\pi n$. Therefore, the value of flow from $A-A^{\prime}$ to $B-B^{\prime}$ is at least $c^{\prime}\pi n-2\cdot\tfrac{1}{3}c^{\prime}\pi n=\tfrac{1}{3}c^{\prime}\pi n$. For each edge $(x,y)\in M_{\tt all}$ with $x\in A-A^{\prime}$, $y\in B-B^{\prime}$ we have $||x-y||^{2}>\Delta$ (otherwise $M$ would not be a maximal matching in $M_{\tt short}$). Therefore,
But then the algorithm should have terminated at line 7 - a contradiction. ∎
@@ 3.3 Matching covers
(c) ${\tt Pr}_{u}[u\mbox{ is regular\,}]\geq 1-\tfrac{1}{n}$.
@@ 3.4 Chaining algorithms
sample $(u_{1},\ldots,u_{K})\sim{\cal N}_{\omega}^{K}|u_{1}$ where $\omega=1-1/K$
  // equivalently, sample $u_{2}\!\sim_{\omega}\!u_{1}$, $u_{3}\!\sim_{\omega}\!u_{2},~{}\ldots,~{}u_{K}\!\sim_{\omega}\!u_{K-1}$
@@
Let $x=\mathbb{E}_{u\sim{\cal N}|u\mbox{\scriptsize~{}is regular},M\sim{\cal M}(u)}[|M|]=\mathbb{E}_{u\sim{\cal N}|u\mbox{\scriptsize~{}is regular},M\sim{\cal M}(u)}[|M^{\tt violating}|]$ (by Lemma 8(b)), $y=\mathbb{E}_{u\sim{\cal N}|u\mbox{\scriptsize~{}is not regular},M\sim{\cal M}(u)}[|M|]\leq n$ and $p={\tt Pr}_{u}[u\mbox{ is not regular}]\leq\tfrac{1}{n}$ (by Lemma 8(c)). We have $n^{1-\varepsilon}\leq\mathbb{E}_{M\sim{\cal M}({\cal N})}[|M|]=(1-p)x+py\leq(1-p)x+\tfrac{1}{n}\cdot n$ and so $(1-p)x\geq n^{1-\varepsilon}-1$. Therefore, $\mathbb{E}_{M\sim{\tt SamplePaths}^{K}({\cal N})}[|M^{\tt violating}|]\geq(1-p)x\geq n^{1-\varepsilon}-1$.